The dimensions of aluminum foil in a box for sale in supermarket are 66 2/3 years by 12 inches. The mass of the foil is 0.83 kg. if its density is 2.70 g/cm3, then what us the thickness of the foil in inches?


1
Expert's answer
2021-02-01T04:42:41-0500

The lengths of the foil in cm is:

l (foil) = 66 2/3 yards = 6096 cm

The width of the foil in cm is:

w (foil) = 12 inches = 30.48 cm

According to the equation:

V = l × w × t

where l - length, w - width, t - thickness

From here:

t = V / (l × w)

According to the equation:

V = m / d

where m - mass, d - density

From here:

t (foil) = m / (d × l × w) = 0.83 kg / (2.70 g/ cm3 × 6096 cm × 30.448 cm) = 830 g / (2.70 g/ cm3 × 6096 cm × 30.448 cm) = 0.001656 cm = 0.00065 inches


Answer: 0.00065 inches